# The Recurrence spectrum for dynamical systems beyond specification

### v1

@@ 2.4. Construction of subsets of the recurrence set
(2.14) $\frac{\log\tau_{\xi_{|\theta^{p}(x)|}(y)}(y)}{|\theta^{p}(x)|+k+t}\leq\frac{\log\tau_{\xi_{n}(y)}(y)}{n}\leq\frac{\log\tau_{\xi_{|\theta^{p}(x)|}(y)}(y)}{|\theta^{p}(x)|}.$
@@
(2.15) $\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y)}(y)}{|\theta^{p+1}(x)|}\leq\frac{\log\tau_{\xi_{n}(y)}(y)}{n}\leq\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y)}(y)}{|\theta^{p}(x)|}.$
@@
$\liminf_{n\to\infty}\frac{\log\tau_{\xi_{n}(y)}(y)}{n}=a\ \text{ and }\ \limsup_{n\to\infty}\frac{\log\tau_{\xi_{n}(y)}(y)}{n}=b,$
namely $y\in R_{\sigma,\xi}(a,b)$. Since $y=y(x)$ and $x\in F_{k}$ is arbitrary, we obtain $\Lambda_{k}(a,b)\subset R_{\sigma,\xi}(a,b)$ as required. ∎
@@ 2.5. Construction of Moran fractals
We continue the same setting as in §2.4.
@@
Since $R(q,\cdot)$ is locally constant on $F_{k}$ for each $q\in\mathbb{N}$ and $\lim_{q\to\infty}R(q,x)=\infty$ for any $x\in F_{k}$, and since $F_{k}$ is compact, we have
(2.16) $\lim_{q\to\infty}\inf_{x\in F_{k}}R(q,x)=\infty.$
Let $x\in F_{k}$. The definition of $y_{q+1}(x)$ in §2.4 implies
@@
$\begin{split}\sum_{p\in R_{q}(x)}|w^{p,1}\theta^{p}(x)w^{p,2}\lambda^{p}(x)w^{p,3}|&\leq\varepsilon e^{\gamma_{\max R(q,x)}\ell_{\max R(q,x)}}\\ &\leq\varepsilon|v^{*}w^{1}(x)v^{1}(x)\cdots w^{M_{\max R(q,x)}(x)}(x)v^{M_{\max R(q,x)}(x)}(x)|\\ &\leq\varepsilon(k+t)(q+2).\end{split}$
@@
From the construction of the seed set $F_{k}$ in §2.4, we have
@@
$\widetilde{\Lambda}_{k}(a,b)=\bigcap_{q=1}^{\infty}\bigcup_{V\in\widetilde{\Delta}_{q}(a,b)}[V],$
where $[V]$ denotes the cylinder in $\Sigma$ spanned by the word $V$. Since $\Lambda_{k}(a,b)=\bigcap_{q=1}^{\infty}\bigcup_{V\in\Delta_{q}(a,b)}[V]$, we have $\widetilde{\Lambda}_{k}(a,b)\subset\Lambda_{k}(a,b)$.
@@
$\inf_{y\in F_{k}}q_{n}(y)\geq\frac{1}{1+2\varepsilon}\frac{n}{k+t}.$
@@
Since $q_{n}$ is locally constant on $\widetilde{\Lambda}_{k}(a,b)$ for each $n\in\mathbb{N}$ and $\lim_{n\to\infty}q_{n}(y)=\infty$ for all $y\in\widetilde{\Lambda}_{k}(a,b)$, and since $\widetilde{\Lambda}_{k}(a,b)$ is compact, we have
(2.20) $\lim_{n\to\infty}\inf_{y\in\widetilde{\Lambda}_{k}(a,b)}q_{n}(y)=\infty.$
By virtue of (2.20) and Lemma 2.10, for all sufficiently large $n\geq 1$ we have $n\leq|y_{q_{n}(y)+1}(f_{a,b}(y))|\leq(1+\varepsilon)(k+t)(q_{n}(y)+2)$ for all $y\in\widetilde{\Lambda}_{k}(a,b)$, and so
$q_{n}(y)\geq\frac{1}{1+\varepsilon}\frac{n}{k+t}-2\geq\frac{1}{1+2\varepsilon}\frac{n}{k+t}.$
@@
We are almost ready to complete the proof of Theorem 1.1. The map $x\in F_{k}\mapsto y(x)\in\Lambda_{k}(a,b)$ is injective. Let $f_{a,b}\colon\Lambda_{k}(a,b)\to F_{k}$ denote the inverse of this map.
@@
Let $V$ belong to the first set in the statement. By the minimality of $q_{n}(y)$, there exists $z\in\widetilde{\Lambda}_{k}(a,b)$ such that $V=y_{q_{n}(y)}(x)wv^{q+1}(x)$ where $x=f_{a,b}(z)$, $y_{q_{n}(y)}(x)\vartriangleleft y_{1}\cdots y_{n}$, $v^{q+1}(x)\in\mathcal{Q}_{k}$, $w\in\mathcal{L}(\Sigma)$ and $|w|\leq t$. The total number of words of this form is bounded from above by $\bigcup_{n=0}^{t}\mathcal{L}_{n}(\Sigma)\cdot\#\mathcal{Q}_{k}\leq m^{k+t+1}$.
Let $V$ belong to the second set in the statement. By the minimality of $q_{n}(y)$, there exists $z\in\widetilde{\Lambda}_{k}(a,b)$ such that $V=y_{q_{n}(y)}(x)w^{p,1}\theta^{p}(x)w^{p,2}\lambda^{p}(x)w^{p,3}v^{q_{n}(y)+1}(x)$, where $x=f_{a,b}(z)$, $y_{q_{n}(y)}(x)\vartriangleleft y_{1}\cdots y_{n}$, $\theta^{p}(x)$, $v^{q_{n}(y)+1}(x)$, $\lambda^{p}(x)\in\mathcal{Q}_{k}$, $w^{p,i}\in\mathcal{L}(\Sigma)$ and $|w^{p,i}|\leq t$ for $i=1,2,3$. Combining Lemma 2.4(c), (2.9) and (2.11) we get
@@
Let $\varphi\colon\Sigma\to\mathbb{R}$ be a constant function. Let $\varepsilon\in(0,h_{\mu}(\sigma))$. Let $k>3t$ be a sufficiently large integer, let $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ be given by Proposition 2.5. Let $a,b\in[0,\infty]$ satisfy $a\leq b$.
@@ 3. Dimension of recurrence sets for interval maps
This section is devoted to the analysis of recurrence sets for interval maps in Theorem 1.2. In §3.1 we introduce a Markov diagram, a directed graph associated with a given interval map. Using a Markov diagram, we show that the coding space of any interval map in Theorem 1.2 satisfies the assumption in Theorem 1.1. In §3.2, we exploit the Markov diagram and the Moran fractals constructed in §2.5 to complete the proof of Theorem 1.2. In §A.2 we prove a statement on the transitivity of the alpha-beta transformations.
@@ 3.1. Markov diagram
Since $T$ is expanding of class $C^{2}$, the function $x\in\pi(X_{T})\mapsto\log|T^{\prime}(\pi^{-1}(x))|\in\mathbb{R}$ is uniformly continuous. Let $\varphi$ denote the continuous extension of this function to $\Sigma_{T}$. Let $k>3t$ be a sufficiently large integer, let $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$, $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ be given by Proposition 2.5 applied to the above decomposition of $\mathcal{L}(\Sigma_{T})$. Let ${\rm diam}(\cdot)$ denote the Euclidean diameter of a subset of $[0,1]$, and set
$K=\inf\{{\rm diam}(\pi^{-1}(D))\colon D\in\mathcal{D}_{N}\}.$
Since $\mathcal{D}_{N}$ is a finite set, $K$ is positive.
Let $a,b\in[0,\infty]$ satisfy $a\leq b$. We have $\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))\subset R_{T,\xi_{T}}(a,b)$. We estimate the Hausdorff dimension of $\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))$ from below. For each $q\in\mathbb{N}$ put
@@
For all sufficiently large integer $k>3t$, all $q\in\mathbb{N}$ and all $B\in\Gamma_{q}(a,b)$, we have
${\rm diam}(B)\geq K\exp(-(\chi_{\nu}(T)+2\varepsilon)kq).$
@@
where the last inequality holds for all sufficiently large $k$. Combining (3.3) and (3.4) yields the desired inequality. ∎
Similarly to §2.6, for each $B\in\Gamma_{q}(a,b)$ fix a point $x(B)$ in ${\rm int}(B)\cap\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))$, and define a Borel probability measure $\nu_{q}$ on $[0,1]$ by
@@
Decreasing $\varepsilon$ to 0 and using (3.2), we obtain the desired equality in Theorem 1.2. The proof of Theorem 1.2 is complete. ∎
@@ A.2. Transitivity of the alpha-beta transformations
Let $S$ denote the set of discontinuities of $T$. Let $U$, $V$ be non-empty open subsets of $[0,1]$. We construct open subintervals $A_{0},A_{1},\ldots$ of $(0,1)$ inductively as follows. Let $A_{0}$ be an arbitrary open interval that is contained in $U$ and does not contain two elements of $S$. Suppose we are given $A_{n}$ for some $n\geq 0$ that does not contain two elements of $S$. If $A_{n}$ is disjoint from $S$, then we set $A_{n+1}=T(A_{n})$. Otherwise, we split $A_{n}\setminus S=J\bigsqcup K$ so that ${\rm diam}(J)>{\rm diam}(K)$, or else ${\rm diam}(J)={\rm diam}(K)$ and $\inf K\geq\sup J$, and set $A_{n+1}=T(K)$. Clearly, if $A_{n}$ and $A_{n+1}$ are defined then ${\rm diam}(A_{n+1})\geq(\beta/2){\rm diam}(A_{n})$. If moreover $A_{n+1}=T(A_{n})$ then ${\rm diam}(A_{n+1})=\beta\cdot{\rm diam}(T(A_{n}))$.
@@
Since $T(I_{0})\cup T(I_{2})=[0,1)$, if (i) or (ii) occurs then we are done. If (iii) occurs, then there is $n^{\prime}\geq n+1$ such that $A_{0}$ contains an open interval that is mapped by $T^{n^{\prime}}$ diffeomorphically onto $(0,1)$. We are done. The proof of Proposition A.1 is complete. ∎
@@ Acknowledgments
Part of this paper was written while I was staying in University of Maryland. I thank James A. Yorke for his hospitality during this visit. This research was partially supported by the JSPS KAKENHI 25K21999, Grant-in-Aid for Challenging Research (Exploratory).

### v2

@@ 2.4. Construction of subsets of the recurrence set
(2.14) $\frac{\log\tau_{\xi_{|\theta^{p}(x)|}(y(x))}(y(x))}{|\theta^{p}(x)|+k+t}\leq\frac{\log\tau_{\xi_{n}(y(x))}(y(x))}{n}\leq\frac{\log\tau_{\xi_{|\theta^{p}(x)|}(y(x))}(y(x))}{|\theta^{p}(x)|}.$
@@
(2.15) $\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y(x))}(y(x))}{|\theta^{p+1}(x)|}\leq\frac{\log\tau_{\xi_{n}(y(x))}(y(x))}{n}\leq\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y(x))}(y(x))}{|\theta^{p}(x)|}.$
@@
$\liminf_{n\to\infty}\frac{\log\tau_{\xi_{n}(y(x))}(y(x))}{n}=a\ \text{ and }\ \limsup_{n\to\infty}\frac{\log\tau_{\xi_{n}(y(x))}(y(x))}{n}=b,$
namely $y(x)\in R_{\sigma,\xi}(a,b)$. Since $x\in F_{k}$ is arbitrary, we obtain $\Lambda_{k}(a,b)\subset R_{\sigma,\xi}(a,b)$ as required. ∎
@@ 2.5. Construction of Moran fractals
We continue the same setting as in section 2.4.
@@
Clearly we have
(2.16) $\lim_{q\to\infty}\max\{R(q,x)\colon x\in F_{k}\}=\infty.$
Let $x\in F_{k}$. The definition of $y_{q+1}(x)$ in section 2.4 implies
@@
$\begin{split}\sum_{p\in R(q,x)}|w^{p,1}\theta^{p}(x)w^{p,2}\lambda^{p}(x)w^{p,3}|&\leq\varepsilon e^{\gamma_{\max R(q,x)}\ell_{\max R(q,x)}}\\ &\leq\varepsilon|v^{*}w^{1}(x)v^{1}(x)\cdots w^{M_{\max R(q,x)}(x)}(x)v^{M_{\max R(q,x)}(x)}(x)|\\ &\leq\varepsilon(k+t)(q+2).\end{split}$
@@
From the construction of the seed set $F_{k}$ in section 2.4, we have
@@
$\widetilde{\Lambda}_{k}(a,b)=\bigcap_{q=1}^{\infty}\bigcup_{V\in\widetilde{\Delta}_{q}(a,b)}[V].$
Since $\Lambda_{k}(a,b)=\bigcap_{q=1}^{\infty}\bigcup_{V\in\Delta_{q}(a,b)}[V]$, we have $\widetilde{\Lambda}_{k}(a,b)\subset\Lambda_{k}(a,b)$.
@@
$\inf\{q_{n}(y)\colon y\in\widetilde{\Lambda}_{k}(a,b)\}\geq\frac{1}{1+2\varepsilon}\frac{n}{k+t}.$
@@
Clearly we have
(2.20) $\lim_{n\to\infty}\inf\{q_{n}(y)\colon y\in\widetilde{\Lambda}_{k}(a,b)\}=\infty.$
By virtue of (2.20) and lemma 2.10, for all sufficiently large $n\geq 1$ we have $n\leq|y_{q_{n}(y(x))+1}(x)|\leq(1+\varepsilon)(k+t)(q_{n}(y(x))+2)$ for all $x\in F_{k}$ with $y(x)\in\tilde{\Lambda}_{k}(a,b)$, and so
$q_{n}(y(x))\geq\frac{1}{1+\varepsilon}\frac{n}{k+t}-2\geq\frac{1}{1+2\varepsilon}\frac{n}{k+t}.$
@@
Notice that the map $x\in F_{k}\mapsto y(x)\in\Lambda_{k}(a,b)$ is injective. Let $f_{a,b}\colon\Lambda_{k}(a,b)\to F_{k}$ denote the inverse of this map.
@@
Let $V$ belong to the first set in the statement of lemma 2.12. By the minimality of $q_{n}(y)$, there exists $z\in\widetilde{\Lambda}_{k}(a,b)$ such that $V=y_{q_{n}(y)}(x)wv^{q+1}(x)$ where $x=f_{a,b}(z)$, $y_{q_{n}(y)}(x)\vartriangleleft y_{1}\cdots y_{n}$, $v^{q+1}(x)\in\mathcal{Q}_{k}$, $w\in\mathcal{L}(\Sigma)$ and $|w|\leq t$. The total number of words of this form is bounded from above by $\bigcup_{n=0}^{t}\mathcal{L}_{n}(\Sigma)\cdot\#\mathcal{Q}_{k}\leq m^{k+t+1}$.
Let $V$ belong to the second set in the statement of lemma 2.12. By the minimality of $q_{n}(y)$, there exists $z\in\widetilde{\Lambda}_{k}(a,b)$ such that
$V=y_{q_{n}(y)}(x)w^{p,1}\theta^{p}(x)w^{p,2}\lambda^{p}(x)w^{p,3}v^{q_{n}(y)+1}(x),$
where $x=f_{a,b}(z)$, $y_{q_{n}(y)}(x)\vartriangleleft y_{1}\cdots y_{n}$, $\theta^{p}(x)$, $v^{q_{n}(y)+1}(x)$, $\lambda^{p}(x)\in\mathcal{Q}_{k}$, $w^{p,i}\in\mathcal{L}(\Sigma)$ and $|w^{p,i}|\leq t$ for $i=1,2,3$. Combining lemma 2.4(c), (2.9) and (2.11) we get
@@
Let $\varphi\colon\Sigma\to\mathbb{R}$ be a constant function. Let $k>\max\{3t,1\}$ be a sufficiently large integer for which there exist $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ as in proposition 2.5. Let $a,b\in[0,\infty]$ satisfy $a\leq b$.
@@ 3. Dimension of recurrence sets for interval maps
This section is devoted to the analysis of recurrence sets for interval maps in theorem 1.2. In section 3.1 we introduce a Markov diagram, a directed graph associated with a given interval map. Using a Markov diagram, we show that the coding space of any interval map in theorem 1.2 satisfies the assumption in theorem 1.1. In section 3.2, we exploit the Markov diagram and the Moran fractals constructed in section 2.5 to complete the proof of theorem 1.2.
@@ 3.1. Markov diagram
Since $T$ is expanding of class $C^{2}$, the function $x\in\pi(X_{T})\mapsto\log|T^{\prime}(\pi^{-1}(x))|\in\mathbb{R}$ is uniformly continuous. Let $\varphi$ denote the continuous extension of this function to $\Sigma_{T}$. Let $k>\max\{3t,1\}$ be a sufficiently large integer for which there exist $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ as in proposition 2.5 applied to the above decomposition of $\mathcal{L}(\Sigma_{T})$. Let ${\rm diam}(\cdot)$ denote the Euclidean diameter of a subset of $[0,1]$, and set
$K_{N}=\inf\{{\rm diam}(\pi^{-1}(D))\colon D\in\mathcal{D}_{N}\}.$
Since $\mathcal{D}_{N}$ is a finite set, $K_{N}$ is positive.
Let $a,b\in[0,\infty]$ satisfy $a\leq b$. We have $\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))\subset R_{T,\xi_{T}}(a,b)$. To estimate the Hausdorff dimension of $\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))$ from below, for each $q\in\mathbb{N}$ we put
@@
For all $q\in\mathbb{N}$ and all $B\in\Gamma_{q}(a,b)$, we have
${\rm diam}(B)\geq K_{N}\exp(-(\chi_{\nu}(T)+2\varepsilon)kq).$
@@
where the last inequality holds provided $k$ is sufficiently large. Combining (3.3) and (3.4) yields the desired inequality. ∎
Similarly to section 2.6, for each $B\in\Gamma_{q}(a,b)$ fix a point $x(B)$ in ${\rm int}(B)\cap\pi^{-1}(\widetilde{\Lambda}_{k}(a,b))$, and define a Borel probability measure $\nu_{q}$ on $[0,1]$ by
@@
Decreasing $\varepsilon$ to 0 and then using (3.2), we obtain the desired equality in theorem 1.2. The proof of theorem 1.2 is complete. ∎
@@ A.2. Transitivity of the alpha-beta transformations
Let $S$ denote the set of discontinuities of $T$. Let $U$, $V$ be non-empty open subsets of $[0,1]$. We construct non-empty open subintervals $A_{0},A_{1},\ldots$ of $(0,1)$ inductively as follows. Let $A_{0}$ be an arbitrary open interval that is contained in $U$ and does not contain two elements of $S$. Suppose we are given $A_{n}$ for some $n\geq 0$ that does not contain two elements of $S$. If $A_{n}$ is disjoint from $S$, then we set $A_{n+1}=T(A_{n})$. Otherwise, we split $A_{n}\setminus S=J\bigsqcup K$ so that ${\rm diam}(J)>{\rm diam}(K)$, or else ${\rm diam}(J)={\rm diam}(K)$ and $\inf K\geq\sup J$, and set $A_{n+1}=T(K)$. Clearly, if $A_{n}$ and $A_{n+1}$ are defined then ${\rm diam}(A_{n+1})\geq(\beta/2){\rm diam}(A_{n})$. If moreover $A_{n+1}=T(A_{n})$ then ${\rm diam}(A_{n+1})=\beta\cdot{\rm diam}(T(A_{n}))$.
@@
Since $T(I_{0})\cup T(I_{2})=[0,1)$, if (i) or (ii) occurs then we are done. If (iii) occurs, then there is $n^{\prime}\geq n+1$ such that $A_{0}$ contains an open interval that is mapped by $T^{n^{\prime}}$ diffeomorphically onto $(0,1)$. The proof of proposition A.1 is complete. ∎
@@ Acknowledgments
I thank anonymous referees and the editor for their careful readings of the manuscript and giving valuable suggestions. Part of this paper was written while I was staying in University of Maryland. I thank James A. Yorke for his hospitality during this visit. This research was partially supported by the JSPS KAKENHI 25K21999.